OpenStudy (anonymous):

A rocket is launched from the surface of the Earth with an upward acceleration of 4 g. after 10 s, what is the speed of the rocket and to what height has it risen?

OpenStudy (anonymous):

V=Vo+at Where V = Final Velocity Vo=Initial " a-acceleration t-time

OpenStudy (anonymous):

V=0+(4m/s^2)(10s)

OpenStudy (anonymous):

V=40m/s

OpenStudy (anonymous):

s=Vo(t)+(0.5)(a)(t^2)

OpenStudy (anonymous):

s=0+(0.5)(4m/s^2)(10s)^2 s=200m

OpenStudy (anonymous):

@christiancsilverio thanks!